Find the asymptotes of the hyperbola x2 −y2 + 2x + y + 9 = 0.


1
Expert's answer
2020-07-03T13:20:11-0400

x2 - y2 + 2x + y + 9 = 0

x2 + 2x - y2 + y = - 9

-(x2 + 2x)/9 + (y2 - y) /9 = 1

-{(x + 1)2 - 1}/9 +{(y - 1/2)2 - 1/4}/9 =

1 + 1/9 - 1/36

{(y - 1/2)2}/9 - {(x+1)2}/9 =1.

The above is the equation of a hyperbola whose transverseaxis is parallel to the y-axis. Thus, the center of the hyperbola is (-1,1/2). Thus, the equations of the two assymptotes of the hyperbola are:

(i) y=1/2 + x - (-1)

y= x + 3/2

(ii)y=1/2 - x + (-1)

y= -x - 1/2